[figure: Header; Body; Footer; Margin Notes; 8; 7; 1; 3; 10; 9; 11; 2; 4; 5; 6]

1  one inch + \hoffset      2  one inch + \voffset
3  \oddsidemargin = 0pt     4  \topmargin = -43pt
5  \headheight = 12pt       6  \headsep = 25pt
7  \textheight = 708pt      8  \textwidth = 469pt
9  \marginparsep = 10pt    10  \marginparwidth = 85pt
11  \footskip = 30pt            \marginparpush = 7pt (not shown)
   \hoffset = 0pt              \voffset = 0pt
   \paperwidth = 597pt         \paperheight = 845pt

[figure: Header; Body; Footer; Margin Notes; 8; 7; 1; 3; 10; 9; 11; 2; 4; 5; 6]

1  one inch + \hoffset         2  one inch + \voffset
3  \evensidemargin = -18pt     4  \topmargin = -43pt
5  \headheight = 12pt          6  \headsep = 25pt
7  \textheight = 708pt         8  \textwidth = 469pt
9  \marginparsep = 10pt       10  \marginparwidth = 85pt
11  \footskip = 30pt               \marginparpush = 7pt (not shown)
   \hoffset = 0pt                 \voffset = 0pt
   \paperwidth = 597pt            \paperheight = 845pt